Sol Summary – April 10th

Crew 279 Sol Summary Report 10-04-2023
Sol: 8
Summary Title: Make hay while the sun shines
Author’s name: Antoine de Barquin, Commander & Agnès Dekeyser, Crew Executive Officer
Mission Status: Nominal
Sol Activity Summary:
Today was a hot day and we had it in our schedule. That’s why the EVA started an hour earlier than usual. The crew had to get up early to be ready to start at 8am. This allowed us not to be exposed to the sun when it was at its highest
Despite this precaution, taking radiation measurements for Thomas’s experiment around the base was no picnic. The hilly terrain behind the station and the already intense heat slowed down our progress. Finally the EVA was a success and we even had the time to take some drone pictures of the station.
In the meantime, the rest of the crew was preparing some good meals for the day. They tried to improve the bread recipe for the second time. Spoiler alert : it might have worked, they will discover the result at tonight’s dinner. They also tried a new recipe : cinnamon rolls! They were delicious but the dough was too hard. It seems that our martionauts have a new recipe to improve. Finally, they cooked some tuna pasta with tomato sauce with some vegetables that the whole crew ate with great pleasure.
The afternoon was filled with work such as taking pictures of the sun for some, turning some videos for others, or working on their own experiment for others. The time is going by extremely fast and our martionauts still have a lot to do but they all remain very motivated!
Look Ahead Plan:
Tomorrow, the crew will check the bacteria samples that were placed at the top of the North Ridge. They will also take this opportunity to perform some drone mapping as well as making some videos for their video report. The view that they will have of the MDRS will be perfect. The next EVA’s will be planned to continue Thomas and Augustin’s experiments meaning radioactivity measurements and drone mapping. In the MDRS, the crew will also continue their own experiment including data analysis and 3D printing.
Anomalies in work: nun-functional robotic observatory and the battery wires of Opportunity needs to be replaced
Weather: Sunny, high 24°C/75°F, low 2°C/36°F
Crew Physical Status: Nominal
EVA: See above
Reports to be filed: Journalist report, Greenhab report, Operations report, EVA report, EVA request
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site